OpenStudy (anonymous):

how do you solve 2x^2+72=0

hartnn (hartnn):

subtract 72 from both sides, what u get ?

hartnn (hartnn):

did u try ?

OpenStudy (anonymous):

its solving a quadratic equation with imaginary solutions.

hartnn (hartnn):

yes, there are quite a few ways to do that, like there is a quadratic formula, do you wish t use that ?

hartnn (hartnn):

the way i was making you solve was : 2x^2=-72 x^2 = -36 taking square root on both sides will give you 2 imaginary roots...

OpenStudy (anonymous):

what i got was x = (plus minus) 6i

hartnn (hartnn):

and you're correct, after taking square root, you get \(x=\pm\sqrt{-36}=\pm \sqrt {-1}\sqrt{36}=\pm 6i\)